40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Jan 4 02:58:53 2011, Jenne, Update, IOO, Power into Mode Cleaner increased 
    Reply  Tue Jan 4 11:06:32 2011, Koji, Update, IOO, Power into Mode Cleaner increased 
       Reply  Tue Jan 4 13:57:48 2011, kiwamu, Update, IOO, incident mirror changed from 45deg to 0 deg 
          Reply  Wed Jan 5 03:06:02 2011, kiwamu, Update, IOO, beam spots on MC mirrors 
       Reply  Tue Jan 4 18:37:18 2011, Jenne, Update, IOO, MCWFS aligned 
Message ID: 4104     Entry time: Tue Jan 4 11:06:32 2011     In reply to: 4103     Reply to this: 4105   4107
Author: Koji 
Type: Update 
Category: IOO 
Subject: Power into Mode Cleaner increased 

- Previously MC TRANS was 9000~10000 when the alignment was good. This means that the MC TRANS PD is saturated if the full power is given.
==> Transimpedance must be changed again.

- Y1-45S has 4% of transmission. Definitively we like to use Y1-0 or anything else. There must be the replaced mirror.
I think Suresh replaced it. So he must remember wher it is.

- We must confirm the beam pointing on the MC mirrors with A2L.

- We must check the MCWFS path alignment and configuration.

- We should take the picture of the new PSL setup in order to update the photo on wiki.

Quote:

What was the point:

I twiddled with several different things this evening to increase the power into the Mode Cleaner.  The goal was to have enough power to be able to see the arm cavity flashes on the CCD cameras, since it's going to be a total pain to lock the IFO if we can't see what the mode structure looks like.

Summed-up list of what I did:

* Found the MC nicely aligned.  Did not ever adjust the MC suspensions.

* Optimized MC Refl DC, using the old "DMM hooked up to DC out" method.

* Removed the temporary BS1-1064-33-1025-45S that was in the MC refl path, and replaced it with the old BS1-1064-IF-2037-C-45S that used to be there.  This undoes the temporary change from elog 3878.  Note however, that Yuta's elog 3892 says that the original mirror was a 1%, not 10% as the sticker indicates. The temporary mirror was in place to get enough light to MC Refl while the laser power was low, but now we don't want to fry the PD.

* Noticed that the MCWFS path is totally wrong.  Someone (Yuta?) wanted to use the MCWFS as a reference, but the steering mirror in front of WFS1 was switched out, and now no beam goes to WFS2 (it's blocked by part of the mount of the new mirror). I have not yet fixed this, since I wasn't using the WFS tonight, and had other things to get done.  We will need to fix this.

* Realigned the MC Refl path to optimize MC Refl again, with the new mirror.

* Replaced the last steering mirror on the PSL table before the beam goes into the chamber from a BS1-1064-33-1025-45S to a Y1-45S.  I would have liked a Y1-0deg mirror, since the angle is closer to 0 than 45, but I couldn't find one.  According to Mott's elog 2392 the CVI Y1-45S is pretty much equally good all the way down to 0deg, so I went with it.  This undoes the change of keeping the laser power in the chambers to a nice safe ~50mW max while we were at atmosphere.

* Put the HWP in front of the laser back to 267deg, from its temporary place of 240deg.  The rotation was to keep the laser power down while we were at atmosphere.  I put the HWP back to the place that Kevin had determined was best in his elog 3818.

* Tried to quickly align the Xarm by touching the BS, ITMX and ETMX.  I might be seeing IR flashes (I blocked the green beam on the ETMX table so I wouldn't be confused.  I unblocked it before finishing for the night) on the CCD for the Xarm, but that might also be wishful thinking.  There's definitely something lighting up / flashing in the ~center of ETMX on the camera, but I can't decide if it's scatter off of a part of the suspension tower, or if it's really the resonance. 

What happened in the end:

The MC Trans signal on the MC Lock screen went up by almost an order of magnitude (from ~3500 to ~32,000).  When the count was near ~20,000 I could barely see the spot on a card, so I'm not worried about the QPD.  I do wonder, however, if we are saturating the ADC. Suresh changed the transimpedance of the MC Trans QPD a while ago (Suresh's elog 3882), and maybe that was a bad idea? 

Xarm not yet locked. 

Can't really see flashes on the Test Mass cameras. 

 

ELOG V3.1.3-